OpenStudy (anonymous):

In which type of relationship will the graph never pass through the coordinate pair (0,0) 1.linear 2.quadratic 3.inverse 4.direct

OpenStudy (anonymous):

None of the above. Examples for each: 1) y=x 2) y=x^2 3) y=-x 4) y=x is a direct relationship as well as linear

OpenStudy (anonymous):

hhmmm -isnt y = 1/x an inverse relationship?

OpenStudy (anonymous):

It is, but he is asking for a class of functions for which (0,0) is never a solution

OpenStudy (anonymous):

Inverse relationship means that as x increases, f(x) decreases. So an example of an inverse relationship for which x=0 is a solution, is f(x)=-x.